solve the proportion. check for extranious solutions.
2/6x+1 = 2x/1
thanks :)

Answer by edjones(7569)

2/(6x+1) = 2x/1 I assume that this is the problem.
2/(6x+1) = 2x
2=2x(6x+1)
12x^2+2x=2
12x^2+2x-2=0
2(6x^2+x-1)=0
6x^2+3x-2x-1=0 6*-1=-6 What factors of -6 add up to +1? Ans: +3 and -2
3x(2x+1)-(2x+1)=0 Factoring by grouping.
(3x-1)(2x+1)=0
.
3x=1
x=1/3
.
2x=-1
x=-1/2
.
Check:
2/(6x+1) = 2x
2/(2+1)=2/3 true
.
2/(6x+1) = 2x
2/(-3+1)=-1 true
.
Ed